# Understanding Representation Dynamics of Diffusion Models via Low-Dimensional Modeling

### v2

@@ 4 Unimodal Representation Dynamics Predicts Model Generalization
Answering this question sheds light on the distribution learning capabilities of diffusion models and is closely related to recent studies on their generalizability, which reveal that diffusion models operate in two distinct regimes—generalization and memorization—depending on factors such as dataset size, model capacity, and training duration (Zhang et al., 2023; Wang et al., 2024; Li et al., 2023b, 2024a; Baptista et al., 2025). In the generalization regime, the model captures the underlying data distribution and generates diverse, novel samples. In contrast, in the memorization regime, it overfits to the training data and loses the ability to generate novel samples.
In this section, using classification tasks as a case study, we empirically demonstrate that the presence of unimodal representation dynamics reliably indicates generalization, while its gradual shift to a monotonically decreasing trend indicates memorization. Specifically, we study the effects of data size and model capacity in Section 4.1, and the effects of learning dynamics in Section 4.2.
@@ 4.1 Effects of dataset size and model capacity on representation dynamics
Recent studies (Zhang et al., 2023; Kadkhodaie et al., 2024) have shown that diffusion models exhibit a phase transition from memorization to generalization as the number of training samples increases. Specifically, when the network size significantly exceeds the number of training samples, diffusion models tend to memorize rather than capture the underlying low-dimensional data distribution (Zhang et al., 2023; Wang et al., 2024). For fixed model capacity, generalization typically emerges when the number of training data is larger than a certain threshold; see Figure 6 (b). Here, we demonstrate that the representation dynamics undergo a similar transition with data scaling. Specifically, we train UNet-32 and UNet-64 (Ronneberger et al., 2015) models using varying training dataset sizes to examine how their representation dynamics change across regimes. The generalization score metric introduced in (Pizzi et al., 2022; Zhang et al., 2023) is used as a reference for quantifying model generalization.
As shown in Figure 6, we observe that reducing the training dataset size leads to a decrease in the generalization score. Correspondingly, the unimodal representation dynamic becomes less obvious, eventually transitioning into a monotonically decreasing curve (i.e., monotonic representation dynamics). These observations highlight a strong connection between representation and distribution learning in diffusion models—specifically, the emergence of unimodal representation dynamics aligns with the ability of the model to capture the underlying data distribution for achieving good generalization.
@@ 4.2 Effects of learning dynamics on representation dynamics
Second, we investigate how learning dynamics influence representation dynamics and also generalization performance, particularly in the limited data regime (e.g., training size $N=2^{12}$ as shown in Figure 6 (b)). In this regime, recent studies (Li et al., 2023b, 2024a; Baptista et al., 2025) have shown that early stopping can improve generalization performance. Specifically, these works observe an early learning phenomenon, where generalization improves during the initial phase of training but deteriorates as the model begins to memorize. As illustrated in Figure 7, this effect is reflected in the evolution of the FID score (Heusel et al., 2017), which initially decreases (indicating better generative quality) and then rises as memorization starts. Notably, we find that this trend negatively correlates with the linear probing accuracy of learned representations. This observation implies that representation quality could potentially serve as an early-stopping criterion to prevent memorization in diffusion models trained on limited data without relying on external models.
@@
Moreover, we show that this early learning behavior can be captured by the transition between unimodal and monotonic curves of representation dynamics during training. Experimentally, we demonstrate this by training EDM-based diffusion models on subsets of CIFAR(Krizhevsky et al., 2009) using $N=2^{12}$ training samples. Specifically, from Figure 8, we observe that the evolution of representation dynamics during training can be divided into two distinct phases:
@@
  • Early phase of unimodal representation dynamics. In the early stage of training, the model learns the underlying low-dimensional data distribution and is able to generalize. As predicted by our theoretical analysis in Section 3, representation follows a unimodal dynamic across noise scales. This unimodal dynamic is clearly observed before training iteration $\mathrm{Iter}\leq 7.5M$ in Figure 8 (a). The generalization behavior is further supported by the new outputs of the model as shown in Figure 8 (b), which more closely resemble those from a reference generalized model than the nearest neighbors in the training set. Moreover, the peak representation quality improves steadily during this phase as the model better captures the data distribution.
@@
  • Late phase of monotonic representation dynamics. However, as training progresses toward convergence, the model begins to memorize the training samples, resulting in a reduced ability to capture the underlying data distribution. This transition is obvious in the outputs of the model, which increasingly replicate training examples (see Figure 8 (b) at $\mathrm{Iter}=15M,100M$). During this phase, the unimodal representation dynamics give way to a monotonic representation dynamics, where representation quality consistently degrades as the noise level increases. Furthermore, as shown in Figure 7, the learned features become less informative, and the peak probing accuracy begins to decline.
@@ 5 Discussion
  • Representation alignment for better image generation. Recent work REPA Yu et al. (2024) has demonstrated that aligning diffusion model features with features from pre-trained self-supervised foundation models can enhance training efficiency and improve generation quality. By providing a deeper understanding of the representation dynamics in diffusion models, our findings could further advance such representation alignment techniques, facilitating the development of diffusion models with superior training and generation performance.
Acnkowledgement
We acknowledge funding support from NSF CAREER CCF-2143904, NSF CCF-2212066, NSF CCF- 2212326, NSF IIS 2312842, NSF IIS 2402950, NSF IIS 2312840, NSF IIS 2402952, ONR N00014-22-1-2529, and MICDE Catalyst Grant.
@@ Appendix A Related works
Denoising autoencoders (DAEs) are trained to reconstruct corrupted images to extract semantically meaningful information, which can be applied to various vision [Vincent et al., 2008, 2010] and language downstream tasks [Lewis et al., 2020]. Related to our analysis of the weight-sharing mechanism, several studies have shown that training with a noise scheduler can enhance downstream performance [Chandra and Sharma, 2014, Geras and Sutton, 2015, Zhang and Zhang, 2018]. On the theoretical side, prior works have studied the learning dynamics [Pretorius et al., 2018, Steck, 2020] and optimization landscape [Kunin et al., 2019] through the simplified linear DAE models.
@@
Diffusion-based representation learning [Fuest et al., 2024] has demonstrated significant success in various downstream tasks, including image classification [Xiang et al., 2023, Mukhopadhyay et al., 2023, Deja et al., 2023], segmentation [Baranchuk et al., 2022], correspondence [Tang et al., 2023], and image editing [Shi et al., 2024]. Using diffusion models for data augmentation has also been shown to improve robustness against covariate shift [Sastry et al., 2023]. To further enhance the utility of diffusion features, knowledge distillation [Yang and Wang, 2023, Li et al., 2023a, Stracke et al., 2024, Luo et al., 2024] methods have been proposed, aiming to bypass the computationally expensive grid search for the optimal $t$ in feature extraction and improving downstream performance. Beyond directly using intermediate features from pre-trained diffusion models, research efforts has also explored novel loss functions [Abstreiter et al., 2021, Wang et al., 2023] and network modifications [Hudson et al., 2024, Preechakul et al., 2022] to develop more unified generative and representation learning capabilities within diffusion models. The work [Han et al., 2025] investigates whether diffusion models are capable of learning latent inter-feature dependencies underlying image data. Unlike the aforementioned efforts, our work focuses more on understanding the representation dynamic of diffusion models and its relationship with model generalization.
@@ Appendix B Auxiliary results
B.1 Extended results from Section 4.2
Figure 9: Representation learning and generative performance across training iterations for Oxford-IIIT Pet dataset [Parkhi et al., 2012]. We train a UNet-128 diffusion model on 3680 training samples in Oxford-IIIT Pet dataset, monitoring both representation learning and generative performance as training progresses. A clear phase transition is observed: early in training, the representations exhibit a unimodal pattern, and generated samples resemble those of a generalizing EDM model, with no signs of memorization. As training continues, the unimodal pattern gradually transitions to a monotonically decreasing trend, aligning with the model’s shift toward memorizing the training data. "NN" denotes nearest neighbor in the training dataset.
Figure 10: Representation learning and generative performance across training iterations for TinyImageNet [Le and Yang, 2015]. We train a UNet-128 diffusion model on $N=2^{11}$ training samples in TinyImageNet, monitoring both representation learning and generative performance as training progresses. A clear phase transition is observed: early in training, the representations exhibit a unimodal pattern, and generated samples resemble those of a generalizing EDM model, with no signs of memorization. As training continues, the unimodal pattern gradually transitions to a monotonically decreasing trend, aligning with the model’s shift toward memorizing the training data. "NN" denotes nearest neighbor in the training dataset.
In Section 4.2, we showed that when training on limited data, the representation dynamics undergo a clear phase transition: from a unimodal pattern to a monotonically decreasing trend, where the diffusion model also exhibits a transition from generalization to increasingly memorize the training data. Here, we provide additional empirical evidence supporting this insight.
We train UNet-based diffusion models using the DDPM++ architecture with the EDM configuration [Karras et al., 2022] on the Oxford-IIIT Pet [Parkhi et al., 2012] and TinyImageNet [Le and Yang, 2015] datasets, using training subsets of 3680 and 2048 images, respectively. Throughout training, we monitor both the evolution of representation dynamics and generative outputs. As shown in Figure 9 and Figure 10, consistent with our findings in Section 4.2, we observe that in the early stages of training, the model exhibits a clear unimodal pattern and generalizes well. As training continues, this unimodal structure gradually shifts into a monotonically decaying trend, and the models start to increasingly replicate training examples.
@@ B.2 Disentangling the role of input noise in representation dynamics
Our theoretical analysis in Section 3 attributes the unimodal pattern to a fundamental trade-off between denoising rate and class confidence rate across noise levels. To validate this explanation and disentangle the effect of additive Gaussian noise, we conduct experiments where feature extraction is performed directly on clean inputs $\bm{x}_{0}$ rather than noisy inputs $\bm{x}_{t}$. We show that the unimodal behavior remains clearly observable even in the absence of injected noise, indicating that the dynamics are not solely a consequence of input corruption. In fact, one can verify that $\mathrm{SNR}(\hat{\bm{x}}_{\bm{\theta}}^{\star}(\bm{x}_{0}),t)$ also exhibits a unimodal trend—this can be shown through an analysis analogous to that of Theorem 1 for $\mathrm{SNR}(\hat{\bm{x}}_{\bm{\theta}}^{\star}(\bm{x}_{t}),t)$.
@@
To support this claim, we train EDM (with the VP configuration) [Karras et al., 2022] on CIFAR datasets and classical DDPM [Ho et al., 2020] on CIFAR10 and Flowers-102 [Nilsback and Zisserman, 2008], and evaluate the representation quality using both noisy inputs $\bm{x}_{t}$ and clean inputs $\bm{x}_{0}$, as shown in Figure 11. For Figure 11(a), we select the layer that yields the best accuracy, while for Figure 11(b), we directly use the bottleneck layer to demonstrate that the observed unimodal behavior is not sensitive to layer choice. Across all settings, the unimodal representation dynamics remain clearly visible, reinforcing that additive Gaussian noise is not the sole factor responsible for this phenomenon.
@@
We visualize the posterior estimation results for clean inputs ($\hat{\bm{x}}_{\bm{\theta}}(\bm{x}_{0},t)$) and noisy inputs ($\hat{\bm{x}}_{\bm{\theta}}(\bm{x}_{t},t)$) across varying noise scales $\sigma_{t}$ in Figure 12. In both cases, the posterior outputs undergo a clear fine-to-coarse transition as $\sigma_{t}$ increases. This supports our theoretical claim that as noise grows, class-irrelevant attributes are gradually removed. The peak in representation quality occurs at an intermediate point where class-essential structure is preserved while irrelevant details are suppressed. When $\sigma_{t}$ becomes too large, class confidence rate drops significantly, resulting in poor representations. The additive noise $\sigma_{t}\bm{\epsilon}$ merely accelerates this degradation but is not the root cause.

### v3

@@ 4 Unimodal Representation Dynamics Predicts Model Generalization
Answering this question sheds light on the distribution learning capabilities of diffusion models and is closely related to recent studies on their generalizability, which reveal that diffusion models operate in two distinct regimes—generalization and memorization—depending on factors such as dataset size, model capacity, and training duration (zhang2024emergence; wang2024diffusion; li2023generalization; li2024understanding; baptista2025memorization). In the generalization regime, the model captures the underlying data distribution and generates diverse, novel samples. In contrast, in the memorization regime, it overfits to the training data and loses the ability to generate novel samples. Further discussion on the generalization of diffusion models are provided in Appendix˜A.
In this section, using classification tasks as a case study, we empirically demonstrate that the presence of unimodal representation dynamics reliably indicates generalization, while its gradual shift to a monotonically decreasing trend indicates memorization. Specifically, we study the effects of data size and model capacity in Section˜4.1, and the effects of learning dynamics in Section˜4.2.
@@ 4.1 Effects of dataset size and model capacity on representation dynamics
Recent studies (zhang2024emergence; kadkhodaie2023generalization) have shown that diffusion models exhibit a phase transition from memorization to generalization as the number of training samples increases. Specifically, when the network size significantly exceeds the number of training samples, diffusion models tend to memorize rather than capture the underlying low-dimensional data distribution (zhang2024emergence; wang2024diffusion). For fixed model capacity, generalization typically emerges when the number of training data is larger than a certain threshold; see Figure˜6 (b). Here, we demonstrate that the representation dynamics undergo a similar transition with data scaling. Specifically, we train UNet-32 and UNet-64 (ronneberger2015u) models using varying training dataset sizes to examine how their representation dynamics change across regimes. The generalization score metric introduced in (pizzi2022self; zhang2024emergence) is used as a reference for quantifying model generalization.
As shown in Figure˜6, we observe that reducing the training dataset size leads to a decrease in the generalization score. Correspondingly, the unimodal representation dynamic becomes less obvious, eventually transitioning into a monotonically decreasing curve (i.e., monotonic representation dynamics). These observations highlight a strong connection between representation and distribution learning in diffusion models—specifically, the emergence of unimodal representation dynamics aligns with the ability of the model to capture the underlying data distribution for achieving good generalization.
@@ 4.2 Effects of learning dynamics on representation dynamics
Second, we investigate how learning dynamics influence representation dynamics and also generalization performance, particularly in the limited data regime (e.g., training size $N=2^{12}$ as shown in Figure˜6 (b)). In this regime, recent studies (li2023generalization; li2024understanding; baptista2025memorization) have shown that early stopping can improve generalization performance. Specifically, these works observe an early learning phenomenon, where generalization improves during the initial phase of training but deteriorates as the model begins to memorize. As illustrated in Figure˜7, this effect is reflected in the evolution of the FID score (heusel2017gans), which initially decreases (indicating better generative quality) and then rises as memorization starts. Notably, we find that this trend negatively correlates with the linear probing accuracy of learned representations. This observation implies that representation quality could potentially serve as an early-stopping criterion to prevent memorization in diffusion models trained on limited data without relying on external models.
@@
Moreover, we show that this early learning behavior can be captured by the transition between unimodal and monotonic curves of representation dynamics during training. Experimentally, we demonstrate this by training EDM-based diffusion models on subsets of CIFAR(krizhevsky2009learning) using $N=2^{12}$ training samples. Specifically, from Figure˜8, we observe that the evolution of representation dynamics during training can be divided into two distinct phases:
@@
  • Early phase of unimodal representation dynamics. In the early stage of training, the model learns the underlying low-dimensional data distribution and is able to generalize. As predicted by our theoretical analysis in Section˜3, representation follows a unimodal dynamic across noise scales. This unimodal dynamic is clearly observed before training iteration $\mathrm{Iter}\leq 7.5M$ in Figure˜8 (a). The generalization behavior is further supported by the new outputs of the model as shown in Figure˜8 (b), which more closely resemble those from a reference generalized model than the nearest neighbors in the training set. Moreover, the peak representation quality improves steadily during this phase as the model better captures the data distribution.
@@
  • Late phase of monotonic representation dynamics. However, as training progresses toward convergence, the model begins to memorize the training samples, resulting in a reduced ability to capture the underlying data distribution. This transition is obvious in the outputs of the model, which increasingly replicate training examples (see Figure˜8 (b) at $\mathrm{Iter}=15M,100M$). During this phase, the unimodal representation dynamics give way to a monotonic representation dynamics, where representation quality consistently degrades as the noise level increases. Furthermore, as shown in Figure˜7, the learned features become less informative, and the peak probing accuracy begins to decline.
@@ 5 Discussion
  • Toward broader theoretical foundations. While our analysis captures key aspects of representation dynamics, it relies on simplified data assumptions and model formulations that facilitate tractable derivations. Extending this framework to more realistic data distributions and complex model architectures remains an important direction for future work. Moreover, establishing a rigorous theoretical connection between the representation geometry of diffusion models and their phase transition between generalization and memorization remains an important and promising direction for future work.
@@
  • Representation alignment for better image generation. Recent work REPA yu2024repa has demonstrated that aligning diffusion model features with features from pre-trained self-supervised foundation models can enhance training efficiency and improve generation quality. By providing a deeper understanding of the representation dynamics in diffusion models, our findings could further advance such representation alignment techniques, facilitating the development of diffusion models with superior training and generation performance.
Acknowledgment
We acknowledge funding support from NSF CAREER CCF-2143904, NSF CCF-2212066, NSF CCF- 2212326, NSF IIS 2312842, NSF IIS 2402950, NSF IIS 2312840, NSF IIS 2402952, ONR N00014-22-1-2529, ONR N000142512339, and MICDE Catalyst Grant. We also thank all the anonymous reviewers for their valuable suggestions and fruitful discussions.
The Appendix is organized as follows: in Appendix˜A, we discuss related works; in Appendix˜B, we present auxiliary findings that complement the main discussion; in Appendix˜C, we provide additional complementary experiments; in Appendix˜D, we present the detailed experimental setups for the empirical results in the paper. Lastly, in Appendix˜E, we provide proof details for Section˜3.
@@ Appendix A Related works
Denoising autoencoders (DAEs) are trained to reconstruct corrupted images to extract semantically meaningful information, which can be applied to various vision [vincent2008extracting, vincent2010stacked] and language downstream tasks [lewis2019bart]. Related to our analysis of the weight-sharing mechanism, several studies have shown that training with a noise scheduler can enhance downstream performance [chandra2014adaptive, geras2014scheduled, zhang2018convolutional]. On the theoretical side, prior works have studied the learning dynamics [pretorius2018learning, steck2020autoencoders] and optimization landscape [kunin2019loss] through the simplified linear DAE models.
@@
Diffusion-based representation learning [fuest2024diffusion] has demonstrated significant success in various downstream tasks, including image classification [xiang2023denoising, mukhopadhyay2023diffusion, deja2023learning], segmentation [baranchuk2021label], correspondence [tang2023emergent], and image editing [shi2024dragdiffusion]. Using diffusion models for data augmentation has also been shown to improve robustness against covariate shift [sastry2023diffaug]. To further enhance the utility of diffusion features, knowledge distillation [yang2023diffusion, li2023dreamteacher, stracke2024cleandift, luo2024diffusion] methods have been proposed, aiming to bypass the computationally expensive grid search for the optimal $t$ in feature extraction and improving downstream performance. Beyond directly using intermediate features from pre-trained diffusion models, research efforts has also explored novel loss functions [abstreiter2021diffusion, wang2023infodiffusion] and network modifications [hudson2024soda, preechakul2022diffusion] to develop more unified generative and representation learning capabilities within diffusion models. The work [han2025diffusionmodelslearnhidden] investigates whether diffusion models are capable of learning latent inter-feature dependencies underlying image data. Unlike the aforementioned efforts, our work focuses more on understanding the representation dynamic of diffusion models and its relationship with model generalization.
Generalization of Diffusion Models.
In this paper, we examine the relationship between a unimodal curve in representation-probing accuracy and the generalization of diffusion models, a topic of active interest [kadkhodaie2023generalization, zhang2024emergence, kamb2024analytic]. There has also been notable work on the mechanisms that enable diffusion models to learn the underlying score function from discrete empirical samples [niedoba2025towards, lukoianov2025locality], thereby generating novel in-distribution samples (i.e., generalizing). We propose a representational perspective on the generalization of diffusion models, together with corresponding measures (unimodality in representation dynamics) for evaluating generalization.
The unimodality of representation-probing accuracy across time steps is also related to symmetry breaking [raya2023spontaneous] and early-stopping generalization during sampling [biroli2024dynamical, sclocchi2025phase] when the model learns empirical score functions. This line of work further connects empirical DMs and dense associative memory (AM) networks through their objectives and sampling behavior [ambrogioni2024search, lucibello2024exponential, pham2025memorization], where generalization is interpreted as novel attraction sinks in AM, which is also related to Gaussian and spherical patterns in the data. However, in this paper our focus is on learning a parameterized model with data from a low-dimensional distribution, an assumption also used in [george2025analysis, achilli2024losing] (which links generalization to adapting to low-dimensional data manifolds and learning local geometric components), rather than on sampling.
@@ Appendix B Auxiliary results
B.1 Extended results from Section˜4.2
Figure 9: Representation learning and generative performance across training iterations for Oxford-IIIT Pet dataset [parkhi2012cats]. We train a UNet-128 diffusion model on 3680 training samples in Oxford-IIIT Pet dataset, monitoring both representation learning and generative performance as training progresses. A clear phase transition is observed: early in training, the representations exhibit a unimodal pattern, and generated samples resemble those of a generalizing EDM model, with no signs of memorization. As training continues, the unimodal pattern gradually transitions to a monotonically decreasing trend, aligning with the model’s shift toward memorizing the training data. "NN" denotes nearest neighbor in the training dataset.
Figure 10: Representation learning and generative performance across training iterations for TinyImageNet [le2015tiny]. We train a UNet-128 diffusion model on $N=2^{11}$ training samples in TinyImageNet, monitoring both representation learning and generative performance as training progresses. A clear phase transition is observed: early in training, the representations exhibit a unimodal pattern, and generated samples resemble those of a generalizing EDM model, with no signs of memorization. As training continues, the unimodal pattern gradually transitions to a monotonically decreasing trend, aligning with the model’s shift toward memorizing the training data. "NN" denotes nearest neighbor in the training dataset.
In Section˜4.2, we showed that when training on limited data, the representation dynamics undergo a clear phase transition: from a unimodal pattern to a monotonically decreasing trend, where the diffusion model also exhibits a transition from generalization to increasingly memorize the training data. Here, we provide additional empirical evidence supporting this insight.
We train UNet-based diffusion models using the DDPM++ architecture with the EDM configuration [karras2022elucidating] on the Oxford-IIIT Pet [parkhi2012cats] and TinyImageNet [le2015tiny] datasets, using training subsets of 3680 and 2048 images, respectively. Throughout training, we monitor both the evolution of representation dynamics and generative outputs. As shown in Figure˜9 and Figure˜10, consistent with our findings in Section˜4.2, we observe that in the early stages of training, the model exhibits a clear unimodal pattern and generalizes well. As training continues, this unimodal structure gradually shifts into a monotonically decaying trend, and the models start to increasingly replicate training examples.
@@ B.2 Disentangling the role of input noise in representation dynamics
Our theoretical analysis in Section˜3 attributes the unimodal pattern to a fundamental trade-off between denoising rate and class confidence rate across noise levels. To validate this explanation and disentangle the effect of additive Gaussian noise, we conduct experiments where feature extraction is performed directly on clean inputs $\bm{x}_{0}$ rather than noisy inputs $\bm{x}_{t}$. We show that the unimodal behavior remains clearly observable even in the absence of injected noise, indicating that the dynamics are not solely a consequence of input corruption. In fact, one can verify that $\mathrm{SNR}(\hat{\bm{x}}_{\bm{\theta}}^{\star}(\bm{x}_{0}),t)$ also exhibits a unimodal trend—this can be shown through an analysis analogous to that of Theorem˜1 for $\mathrm{SNR}(\hat{\bm{x}}_{\bm{\theta}}^{\star}(\bm{x}_{t}),t)$.
@@
To support this claim, we train EDM (with the VP configuration) [karras2022elucidating] on CIFAR datasets and classical DDPM [ho2020denoising] on CIFAR10 and Flowers-102 [nilsback2008automated], and evaluate the representation quality using both noisy inputs $\bm{x}_{t}$ and clean inputs $\bm{x}_{0}$, as shown in Figure˜11. For Figure˜11(a), we select the layer that yields the best accuracy, while for Figure˜11(b), we directly use the bottleneck layer to demonstrate that the observed unimodal behavior is not sensitive to layer choice. Across all settings, the unimodal representation dynamics remain clearly visible, reinforcing that additive Gaussian noise is not the sole factor responsible for this phenomenon.
@@
We visualize the posterior estimation results for clean inputs ($\hat{\bm{x}}_{\bm{\theta}}(\bm{x}_{0},t)$) and noisy inputs ($\hat{\bm{x}}_{\bm{\theta}}(\bm{x}_{t},t)$) across varying noise scales $\sigma_{t}$ in Figure˜12. In both cases, the posterior outputs undergo a clear fine-to-coarse transition as $\sigma_{t}$ increases. This supports our theoretical claim that as noise grows, class-irrelevant attributes are gradually removed. The peak in representation quality occurs at an intermediate point where class-essential structure is preserved while irrelevant details are suppressed. When $\sigma_{t}$ becomes too large, class confidence rate drops significantly, resulting in poor representations. The additive noise $\sigma_{t}\bm{\epsilon}$ merely accelerates this degradation but is not the root cause.